OpenStudy (anonymous):

Integral sqrt ( 1 - x^2) dx, from -1 to 1.

OpenStudy (anonymous):

try trig substitution: x=sin(u) dx=cos (u) du let me know if there is a problem

OpenStudy (anonymous):

I sqrt ( 1 - sin^2 u) cos u du = I cos u cos u du = I cos^2 u du And then?

OpenStudy (anonymous):

cos^2x = 1/2 (1+cos2x) and take integral of each term

OpenStudy (anonymous):

I got: = 1/2 u +1/4 sin2u

OpenStudy (anonymous):

Ok. I cos^2 u du = 1/2 I 1 + cos 2 u du 1/2 * u + 1/2 sin 2 u / 2 = 1/2 u + 1/4 sin 2u

OpenStudy (anonymous):

now, what it the u=...? x=sin u then u=arcsin x right?

OpenStudy (anonymous):

sin2u=2sin u * cos u = \[=2x*\sqrt{1-x^{2}}\]

OpenStudy (anonymous):

apply your (-1,1) to the final answer

OpenStudy (anonymous):

The, so I can work straight with u, I'll change - 1 to 1 to sin u = - 1 to sin u = 1 u = 270° and 90°. 1/2 * 3 pi /2 + 1/4 * 0 - ( 1/2 * pi/2 + 1/4 * 0 ) = 3 pi / 4 - 1/4 * pi = 2/4 pi = 1/2 pi

OpenStudy (anonymous):

should work! is your answer matched?

OpenStudy (anonymous):

No wait a minute, sin pi = 1. So: 3 pi /4 - (1/4 pi + 1/4) = 2/4 pi + 1/4....

OpenStudy (anonymous):

2/4 pi - 1/4.

OpenStudy (anonymous):

It's wrong though, the answer should be pi.

OpenStudy (anonymous):

I didn't calculate... should I?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

give me a min

OpenStudy (anonymous):

Ok.

OpenStudy (anonymous):

i got -pi/2....?

OpenStudy (anonymous):

Ah, you know what? Let's just leave it aside for a moment. Thanks for the help, sometime I'll still get round to do it.

OpenStudy (anonymous):

OK. what it the answer should be - I'll play with it :)

OpenStudy (anonymous):

The answer should be pi, see, this integral gives the area of a r=1 circle.

OpenStudy (anonymous):

oooo.... because it's sq rt - graph is symmetrical over y-axis!

OpenStudy (anonymous):

it should be : \[=2*\int\limits_{0}^{\pi/2}\cos ^{2}u du=....\] =2(1/2u+1/4 sin2u)=u + 1/2 sin2u= =pi +0=pi Yes!